OpenStudy (anonymous):

what is the simplified version of: 1/2 log 3 x – 2 log 3 y - 4 log 3 z

OpenStudy (anonymous):

log[(3x)^1/2/((3y)^2*(3z)^4)]

OpenStudy (anonymous):

the choices are 3sqrtxy^2/z^4

OpenStudy (anonymous):

3sqrtxz^4/y^2

OpenStudy (anonymous):

3 sqrt x/y^2 z^4

OpenStudy (anonymous):

\[\log_{10} {\frac{\sqrt{3x}}{9y ^{2}81z ^{4}}}\]

OpenStudy (anonymous):

or 3x^3 yz

OpenStudy (anonymous):

those arent any of the answer choices =/

OpenStudy (anonymous):

\[\frac{1}{2}\log(3x)-2\log(3y)-4\log(3z)\] is this your question

OpenStudy (anonymous):

\[1/2\log_{3} x-2\log_{3} y-4\log_{3} z\] if you asking this \[\log_{3} (\frac { \sqrt{x}}{y ^{2}z ^{4}})\]